A friend of yours bought a new sports car with a $5,500 down payment plus a $32,000 car loan that is financed at an interest rate of 0.75% per month for 60 months. After 2 years, the Blue Book value of her vehicle in the used-car marketplace is $19,000 a. Calculate the required monthly loan payment on the car b. How much does your friend still owe on the car loan immediately after she makes her 24th payment? d. If she decides to keep the car but pay $300 more each month, how many months will it take her to payoff the remaining loan she owes? a. The required monthly payment is s(Round to the nearest cent.) /c. Compare your answer to Part (b) to $19,000. This situation is called being upside down. What can she do about it? Your friend still owes Son the car loan. (Round to the nearest dollar.) C.What is the best thing to do in this situation? Choose the correct answer below. A. She should stop paying out her loan as it is obviously unfair O B. She should sell the car immediately O C. In case the mileage is high, the car is worth more than $19,000 and she should consider selling it O D. She should keep the car longer and hope the vehicle remains in good working order d. It will take her months to payoff the remaining loan balance (Round-up to the nearest month) Best
Answer #1

Given that,

Downpayment = 5500

Loan Amount = 32000

Interest Rate = 0.75% per month

Loan Tenure = 60 months

a)

Monthly Payment PMT = (0.75%,60,32000) = 664.27 per month (Using excel formula (Rate,Tenure,PV))

b) Total Payment made in 24 month = 24 * 664*27 = 15942.48

Total Payment to be made in 60 months = 60*664.27 = 39856.2

Payment due after 24 months = 39856.2 - 15942.48 = 23913.72

c) The correct option will a) she should stop paying out her loan as it is unfair to pay the loan more that 19000.